OpenStudy (anonymous):

This is so hard! ): What is f(x)= 2x^2 + 5 Multiplied By G(x)=2x A. 7x+2x B. 4x+10x C. 4x^2+10x D. 4^3+10x

OpenStudy (anonymous):

2x(2x^2+5)= ??

OpenStudy (zzr0ck3r):

\[(2x^2+5)(2x) = (2x)*(2x^2)+(2x)*(5) = 2*2x^{2+1}+10x=4x^3+10x\]

OpenStudy (anonymous):

oh okay thanks for your help @zzr0ck3r youve helped me before! and is the answer 4^3 + 10x?

OpenStudy (zzr0ck3r):

yes

OpenStudy (anonymous):

thanks! your a life saver!

OpenStudy (zzr0ck3r):

all we did was this a(b+c) = ab+ac

OpenStudy (anonymous):

oh okay i see thanksfor the help!